OpenStudy (anonymous):

ALGEBRA ONE HONORS Which of the following will give you the incorrect slope? A--y squared - y to 1 power over --x squared - x to 1 power B--y squared - y to 1 power over --x to 1 power - x squared C--y to 1 power - x squared over --x to 1 power - x squared D--rise over --run

OpenStudy (phi):

You must mean \(y_1 \) and \(y_2 \) not \(y^1\) or \(y^2\) given two points on a line ( \(x_1,y_1\)) and ( \(x_2,y_2\)) the slope if the line is the change in y over the change in x \[ m= \frac{y_1-y_2}{x_1-x_2} \] order of the two points does not change things: \[ m= \frac{-(y_1-y_2)}{-(x_1-x_2)} = \frac{y_2-y_1}{x_2-x_1} \] Slope is also known as rise over run (means the same as the formula) If you look at your choices both B and C are not valid expressions for the slope (I suspect there is a typo in B)

OpenStudy (anonymous):

it should be B) y squared - x to 1 power over -- x to 1 power - x squared

OpenStudy (phi):

in that case C must have a typo

OpenStudy (phi):

and it is not y squared it is y2 (the y value of the 2nd point) y^2 means something entirely different.

OpenStudy (anonymous):

okay so C is the answer?

OpenStudy (phi):

as posted both B and C are not slopes